It has been suggusted that the highest priority of retirees is travel. Thus, a study was conducted to investigate the differences in the length of stay of a trip for pre- and post-retirees. A sample of 700 travelers were asked how long they stayed on a typical trip. The observed results of the study are found below. Number of Nights Pre-retirement Post-retirement Total 4−7 238 172 410 8−13 84 62 146 14−21 33 58 91 22 or more 17 36 53 Total 372 328 700 With this information, construct a table of estimated expected values. Number of Nights Pre-retirement Post-retirement 4−7 8−13 14−21 22 or more Now, with that information, determine whether the length of stay is independent of retirement using α=0.05. (a) χ2= (b) Find the degrees of freedom: (c) Find the critical value: (d) The final conclusion is A. We can reject the null hypothesis that the length of stay is independent of retirement and accept the alternative hypothesis that the two are dependent. B. There is not sufficient evidence to reject the null hypothesis that the length of stay is independent of retirement.
Answer #1

a)

Applying chi square test:

Expected Ei=row total*column total/grand total Pre Post Total
4-7 217.886 192.114 410
8-13 77.589 68.411 146
14-21 48.360 42.640 91
22 or more 28.166 24.834 53
total 372 328 700
chi square    χ2 =(Oi-Ei)2/Ei Pre Post Total
4-7 1.857 2.106 3.963
8-13 0.530 0.601 1.131
14-21 4.879 5.533 10.412
22 or more 4.426 5.020 9.447
total 11.692 13.260 24.952

χ2= 24.952

b)

degree of freedom(df) =(rows-1)*(columns-1)= 3

c)

for 3 df and 0.05 level of signifcance critical value       χ2= 7.815

d)

A. We can reject the null hypothesis that the length of stay is independent of retirement and accept the alternative hypothesis that the two are dependent.